The intrinsic work functions of some elements are listed below. Determine the minimum wavelength/region in the electromagnetic spectrum that can remove the least strongly held electron of that material. Refer to the wavelengths below, if there is none, write none.

Yellow (526x1012Hz), Green (606x1012Hz), Red (484x1012Hz), Violet (789x1012Hz), Orange (508x1012Hz), Blue (668x1012Hz), Near Infrared (214x1012Hz), Mid-Infrared (37x1012Hz), Far-Infrared (0.3x1012Hz), Near Ultraviolet (4.13eV), and Far Ultraviolet (12.4eV).

Silicon – 4.85eV Lithium – 2.0 eV Gold – 5.47 eV
Answer #1

for silicon work function W = 4.85 eV

E = hf

4.85*1.6*10^-19 = 6.63*10^-34*f

f = 1170*10^12 Hz

it is not possible to remove electron from silicon with given electromagnetci spectrum

None


for Lithium = 2eV

2*1.6*10^-19 = 6.63*10^-34*f

f = 483*10^12 Hz

for Lithium, Yellow green , red

for Gold = 5.47 eV

5.47*1.6*10^-19 = 6.63*10^-34*f

f = 1321 Hz

it is not possible to remove electron from gold with given electromagnetci spectrum

None